James has more dimes than quarters. If he has $7.25, how many of each coin does he have?
I have tried 10x + 25x = 725, but it doesn't seem right since he is supposed to have more dimes than quarters.
Thanks!

Answer by ewatrrr(24785)

25a + 10b = 725
a > b
Re TY, noticed(if a = b) 725/35 = 20+, so tried 21 for a
25(21) + 10(20) = 725cents
21 Quarters and 20 Dimes